Suppose that consumer want to spend his all income in goods A and B and the goods are woarthless without one another and the price for both is 3
1. How many units of A and B consumer can buy.
2.if price of A is increase to 6 how many units of A and B can buy.
3.what types of consumption does two types of goods represent
1
Expert's answer
2021-08-12T05:09:06-0400

1)lets take i as the income.

"a=\frac{1}{3a}\\b=\frac{i}{3b}"

2) price of a will be 3+6=9

"a=\frac{1}{9a}\\b=\frac{i}{3b}"

3)The goods are compliments.